# Do assumptions about the central density of subhaloes affect dark matter annihilation and lensing calculations?

### v2

@@ 4.2 $J$-Factor Evolution
Ultimately we are interested in predicting the total annihilation rate in dark matter subhaloes, which is proportional to the $J$-factor defined in Equation 4.2. In Figure 7, we show predictions for how the $J$-factor evolves with mass loss for the different profile models. We calculate the $J$-factor within the tidal radius but note that the $J$-factor is insensitive to this choice. As with the boost factor, we find large differences in the $J$-factor between the different models. By the time the profile is stripped to 5 per cent of the initial mass, there is a factor of 5 difference in the annihilation signal between the ST approximation and the H03 approximation, while the Energy Truncation predictions lie between the two. It is often assumed that since the high-density inner regions of subhaloes are resilient against tidal forces, the annihilation rate will remain fairly constant during tidal stripping [4]. Figure 7 shows that this assumption is incorrect, once mass loss reaches a significant level.
@@
Comparing the three Energy Truncation models, the EinLowT $J$-factor (whose density profile has a steeper central cusp; Figure 2) decreases the least, while the value from the flatter EinHigh profiles decreases the most. The ST model (dashed line) always overestimates the annihilation signal. Although the H03 model (dotted line) gives a reasonable match to the stripped density profile except at very low masses/radii (Figure 3), this model predicts much lower annihilation rates overall.
The variation in the predicted $J$-factor between models reflects the sensitivity of the annihilation signal to the central density. In Figure 8 we show how the enclosed central density evolves with bound mass. We consider the mean density within 10 per cent ($\bar{\rho}_{0.1}$; top panel) and 1 per cent ($\bar{\rho}_{0.01}$; bottom panel) of the initial virial radius (as described in Section 3). The cuspy EinLow profile has the smallest change in central density, which is why the $J$-factor for this model changes slowly. When comparing approximations for tidally stripped NFW profiles, the central density of the ST model is constant as expected, as long as the tidal radius exceeds the radius used to calculate the enclosed density. Relative to the NFW results, both the ST and the H03 model do a fairly good job at predicting the enclosed density within 10 per cent of the virial radius. However, the ST approximation over-predicts the central density within 1 per cent of the virial radius, while the H03 approximation greatly under-predicts the enclosed density, as discussed in Section 2.3.
Figure 8: Change in the central density versus bound mass fraction for various profiles. The enclosed density is calculated in either 10 per cent (top) or 1 per cent (bottom) of the initial virial radius. The enclosed density decreases over time, with slight differences between the three Energy Truncation profiles. For the ST approximation, the enclosed density is constant over time within the tidal radius and thus over-predicts the more accurate NFWT model. The H03 approximation predicts a large decrease in the central density and thus under-predicts the NFWT model.
@@ 4.3 Dependence on structural parameters
For an NFW profile, the boost factor can be well-approximated as a function of the concentration parameter $c=r_{\rm vir}/r_{-2}$ alone, with $B\sim c^{2.5}$, while the boost factor for Einasto profiles scales similarly with a slight dependence on the shape parameter $\alpha$ [68]. For tidally stripped haloes, $r_{t}$ is a more natural outer radius, as discussed in Section 7. In Figure 9, we show the relationship between the boost factor and the effective concentration parameter, $c_{t}$. We also compare this relation to the [68] parameterization for how $B$ varies with $c$ for an untruncated NFW profile (the dash-dot grey line). As mentioned in [68], we expect that $B(c_{t})$ should be approximately equal between the unstripped and stripped profiles since most of the mass is removed outside the tidal radius. We find that though $B(c_{t})$ matches the un-truncated NFW calculation for the H03 and ST models, our three Energy Truncation models have much higher boost factors for a given tidal radius, as shown previously in Figure 6.
@@
Figure 9: Boost factor as a function of the effective concentration parameter, $c_{t}=r_{t}/r_{-2}$. For comparison, we show the expected relation for an unstripped NFW profile (dotted gray line), using the parameterization from okoli2018 [68]. Note that for the Energy-Truncation models, there are some cases where $r_{t}>r_{\rm vir}$, and therefore not a physically accurate depiction of the concentration. We denote cases where $r_{t}>10r_{s}$ with a dashed line. Overall, we find that the Energy Truncation models have boost factors at least twice as large as expected from the unstripped profiles; this is contrary to the expectation that the boost factor of the stripped profile should be approximately equal to that of the unstripped profile within the tidal radius.
@@
The $J$-factor and structural properties of subhaloes are also known to be related. For instance, [14] examined the relationship between $J$ and $G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max}$, where $v_{\rm max}$ and $r_{\rm max}$ are the maximum circular velocity ($v_{c}=\sqrt{GM(<r)/r}$), and corresponding radius. They show that while for an NFW profile $\log_{10}J=1.23\log_{10}(G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max})$, their tidally-stripped simulations of NFW profiles have a slightly shallower slope of 0.86. Our NFWT model agrees remarkably well with their findings, but differences in the profile model can cause a difference in the $J$-factor by an order of $\approx 2$.
@@
Figure 10: Relationship between the $J$-factor, and the combination of structural parameters $G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max}$, where $v_{\rm max}$ and $r_{\rm max}$ are the maximum circular velocity. The slope of 1.23 is what is expected for an NFW profile, while 1.86 is what has been found for tidally stripped NFW profiles [14]. We find that our NFWT model agrees with this scaling relation, but differences in initial profiles can cause a deviation in the $J$-factor by factors of $\approx 2$.
@@ 4.4 Model Sensitivity in Annihilation: Summary
In this section, we considered the effects of tidal mass loss on the inner structure of subhaloes as they relate to calculations of the annihilation rate. Assuming a truncation of the distribution function in energy space, we calculated quantities related to the dark matter annihilation rate for NFW, EinHigh, and EinLow profiles. Additionally, we compared our prediction for NFW profiles to two widely used approximations; first, a model in which the original NFW profile is sharply truncated such that it contains the same bound mass (ST model), and secondly, the parameterization from [39] that was tuned to simulation results (H03).
@@
Comparing first the energy-truncated NFW and Einasto profiles, we found that for fixed mass, the concentration, boost factor (measured within the tidal radius), and $J$-factor of the EinLowT profile is typically 10-50 per cent higher than the NFWT profile, while the EinHighT profile is typically lower. These differences reflect the different central densities of the initial models, as the annihilation signal is dominated by these high-density regions.
@@
Then, comparing tidal stripping models (the NFWT model versus the commonly used ST and H03 approximations), we found that neither of the common approximations agrees well with the Energy Truncation model. The effective concentration is lower by a factor of 10 for both approximations. Both models under-predict the boost factor within the tidal radius by a factor of $\sim 100$. However, this result is mainly due to the differences in the tidal radii, and if we consider the evolution of $B/B_{init}$ the H03 and ST model underestimate and overestimate the boost factor by $\sim 20$ per cent, respectively. If we consider the $B/B_{\rm un}$; i.e., the boost factor from our model compared to a truncated profile within the same radius, $r_{\rm t}$, we can find that the two approximations underestimate the boost factor by an order of magnitude. Overall, the $J$-factor is overestimated in the ST model and underestimated in the H03 model; the differences between these models are a factor of 5 once the halo has been stripped to $\sim 5\%$ of its initial mass.
@@
In summary, subhalo concentrations, boost factors, and $J$-factors of tidally stripped haloes are sensitive to the innermost part of the density profile, and predicted fluxes can vary by a factor of 5 or more depending on profile assumptions. Thus, it is important to have physically motivated models, such as the Energy Truncation model, that allow us to predict the annihilation of subhaloes down to arbitrarily small bound mass fractions. Our results suggest that current constraints on the CDM annihilation cross-section may be inaccurate due to their underlying assumptions about subhalo density profiles; addressing these inaccuracies will be the focus of future work.
@@ 4.5 Implications for the Milky Way
We provide a brief exploration of what our results mean in the context of the Milky Way galaxy — more robust calculations will be the focus of future work. Throughout this section, we assume the Milky Way has a virial mass of $M_{\rm MW}=10^{12}M_{\odot}$, a virial radius of 250 kpc, and a concentration of 10. We use a Planck 2018 cosmology [71].
The total Milky Way annihilation signal can be expressed as
$J_{\rm MW}=J_{\rm sub}+J_{\rm strip}\,\,\,,$ (4.5)
where $J_{\rm sub}$ is the annihilation signal due to Milky Way subhaloes at redshift $z=0$ and $J_{\rm strip}$ is the annihilation signal from the stripped mass.
The signal from the substructure can be found by adding the contribution from each subhalo,
$J_{\rm sub}=\int_{\infty}^{0}\int_{M_{\rm min}}^{M_{\rm max}}\dfrac{{\rm d}N}{{\rm d}M_{i}{\rm d}z}J(M_{0}){\rm d}M_{i}{\rm d}z\,\,\,,$ (4.6)
where $N$ is the number of haloes, $M_{i}$ is the mass of a subhalo at infall redshift, $z$, and $M_{0}$ is the mass of the subhalo at redshift 0. The annihilation signal of each subhalo, $J=J(M_{0})$, depends on the assumed density profile of the subhalo. We assume a minimum halo mass of $10^{-6}\,M_{\odot}$, as in [68].
To calculate $J_{\rm strip}$, we assume that the stripped mass is smoothly distributed in an NFW profile, and is given by:
$M_{\rm strip}=\int_{\infty}^{0}\int_{M_{\rm min}}^{M_{\rm max}}\dfrac{{\rm d}N}{{\rm d}M_{i}{\rm d}z}(M_{i}-M_{0}){\rm d}M_{i}{\rm d}z\,\,\,.$ (4.7)
To find the relationship between the infall mass and the current mass of the subhaloes, we use [48]:
$M_{0}=M_{i}\exp\left(\dfrac{-0.81\sqrt{200}z_{i}}{\pi}\right)\,\,\,,$ (4.8)
where $z_{i}$ is the infall redshift, and we calculate $J_{\rm sub}$ using different models for the subhalo density profiles. However, we note that this parameterization is likely an overestimate of the mass loss, and does not take into account the effect of subhalo structure on mass-loss rates. Finally, we assign concentrations to subhaloes assuming they lie on the median concentration–mass relation from [46].
We calculate ${\rm d}N/{\rm d}M_{i}{\rm d}z$ by assuming the infall population is the same as the field population:
$\dfrac{{\rm d}N}{{\rm d}M_{i}{\rm d}z}=\dfrac{1}{\rho_{c}(z)}\dfrac{{\rm d}M_{\rm MW}(z)}{{\rm d}z}\dfrac{{\rm d}n}{{\rm d}M}\,\,\,,$ (4.9)
@@
where ${\rm d}n(M,z)/{\rm d}M$ is the [20] halo mass function, $M_{\rm MW}(z)$ is the mass of the Milky Way at redshift $z$, and $\rho_{c}$ is the critical density. We assume the Milky Way mass evolves according to [87]:
$M_{\rm MW}(z)=M_{\rm MW}(0)\exp(-8.2z/c)\,\,\,.$ (4.10)
@@
The top panel of Figure 11 shows the relative contribution of substructure to the total annihilation signal, as a function of subhalo mass. The signal is dominated by the smallest subhaloes, with approximately 70 per cent of the signal coming from haloes less than $10^{6}\,M_{\odot}$. The contribution of subhaloes to the total annihilation signal differs slightly between models, with small subhaloes in more centrally concentrated models contributing slightly more to the annihilation signal.
@@
Then, we compared the ‘substructure boost factor’ [68] — defined as the annihilation flux from substructure, $J_{\rm sub}$ relative to the flux from the smooth component of the halo — in the bottom panel Fig 11. We calculated the smooth component by assuming all of the Milky Way mass was smoothly distributed in an NFW profile with mass
@@
We find that including substructure increases the signal by a factor of $\sim 50$ for the more centrally concentrated profiles (ST and EinLowT), and $\sim 10$ times for the less centrally concentrated profiles (H03). The lower values are comparable to the predictions of [68], who explored different concentration–mass relations, but found a substructure boost factor of $\sim$16 when assuming a H03 model and concentration relations similar to ours [^5] We note that our quantity is slightly different — we are showing the substructure annihilation rate to the smooth signal using the total Milky Way Mass, while [68] uses the smooth signal from the mass not contained in subhaloes.. This demonstrates the overall importance of substructure in calculating the annihilation signal in the Milky Way, but also the large range of values predicted depending on the assumed subhalo density profile. Assuming low-mass subhaloes have profiles closest to the EinLowT or NFWT models, we predict a ‘substructure boost factor’ of as much as 20–50, relative to the contribution from the smooth halo.
We note, however, that there are several simplifications in this analysis. First, our parameterization for subhalo stripping (Equation 4.8) overestimates how quickly profiles are disrupted. Second, we assumed that all haloes lie on the median concentration–mass relation and that the Milky Way has built up its mass at an average rate. Finally, the rate at which haloes are stripped is dependent on the subhalo model, with cuspier profiles experiencing less stripping — this, however, would increase the difference between the models. We leave a full investigation of how subhalo evolution affects the annihilation signal for future work.

### v3

@@ 4.2 $J$-factor evolution
Ultimately we are interested in predicting the total annihilation rate in dark matter subhaloes, which is proportional to the $J$-factor defined in Equation 4.2. In Figure 7, we show predictions for how the $J$-factor evolves with mass loss for the different profile models. We calculate the $J$-factor within the tidal radius but note that the $J$-factor is insensitive to this choice. As with the flux multiplier , we find large differences in the $J$-factor between the different models. By the time the profile is stripped to 5 per cent of the initial mass, there is a factor of 5 difference in the annihilation signal between the ST approximation and the H03 approximation, while the Energy Truncation predictions lie between the two. It is often assumed that since the high-density inner regions of subhaloes are resilient against tidal forces, the annihilation rate will remain fairly constant during tidal stripping [59]. Figure 7 shows that this assumption is incorrect, once mass loss reaches a significant level.
@@
Comparing the three Energy Truncation models, the $J$-factor in the EinCuspT case (whose density profile has a steeper central cusp; Figure 2) decreases the least, while the value from the flatter EinCore profiles decreases the most. The ST model (dashed line) is expected to always overestimate the annihilation signal. Although the H03 model (dotted line) gives a reasonable match to the stripped density profile except at very low masses/radii (Figure 3), this model predicts much lower annihilation rates overall.
The variation in the predicted $J$-factor between models reflects the sensitivity of the annihilation signal to the central density. In Figure 8 we show how the enclosed central density evolves with bound mass. We consider the mean density within 10 per cent ($\bar{\rho}_{0.1}$; top panel) and 1 per cent ($\bar{\rho}_{0.01}$; bottom panel) of the initial virial radius (as described in Section 3). The cuspy EinCusp profile has the smallest change in central density, which is why the $J$-factor for this model changes slowly. When comparing approximations for tidally stripped NFW profiles, the central density of the ST model is constant as expected, as long as the tidal radius exceeds the radius used to calculate the enclosed density. Relative to the NFW results, both the ST and the H03 model predictions agree closely with the NFWT model for the enclosed density within 10 per cent of the virial radius. However, the ST approximation over-predicts the central density within 1 per cent of the virial radius, while the H03 predicts a signficantly lower enclosed density, as discussed in Section 2.3.
Figure 8: Change in the central density versus bound mass fraction for various profiles. The enclosed density is calculated in either 10 per cent (top) or 1 per cent (bottom) of the initial virial radius. The enclosed density decreases over time, with slight differences between the three Energy Truncation profiles. For the ST approximation, the enclosed density is constant over time within the tidal radius and thus over-predicts the more accurate NFWT model. The H03 approximation—which is known to underestimate the central density—predicts a significantly larger decrease in the central density compared to the NFW model.
@@ 4.3 Dependence on structural parameters
For an NFW profile, the flux multiplier can be well-approximated as a function of the concentration parameter $c=r_{\rm vir}/r_{-2}$ alone, with $f\sim c^{2.5}$, while the flux multiplier for Einasto profiles scales similarly with a slight dependence on the shape parameter $\alpha$ [2]. However, for tidally stripped haloes, $c_{t}$ is a more natural concentration parameter, as discussed in Section 3. In Figure 9, we show the relationship between the flux multiplier and the effective concentration parameter, $c_{t}$. We also compare this relation to the [2] parameterization for how $f_{M}$ varies with $c$ for an untruncated NFW profile (the dash-dot grey line). As mentioned in [2], we expect that $f_{M}(c_{t})$ should be approximately equal between the unstripped and stripped profiles since most of the mass is removed outside the tidal radius. We find that though $f_{M}(c_{t})$ matches the un-truncated NFW calculation for the H03 and ST models, our three Energy Truncation models have much higher flux multipliers for a given tidal radius, as shown previously in Figure 6.
@@
Figure 9: Flux multiplier as a function of the effective concentration parameter, $c_{t}=r_{t}/r_{-2}$. For comparison, we show the expected relation for an unstripped NFW profile (dashed-dotted gray line), using the parameterization from [2]. Note that for the Energy-Truncation models, there are some cases where $r_{t}>r_{\rm vir}$, and therefore is not a physically accurate depiction of the concentration. We denote cases where $r_{t}>10r_{s}$ with a dashed line. Overall, we find that the Energy Truncation models have flux multipliers at least twice as large as expected from the unstripped profiles; this is contrary to the expectation that the flux multiplier of the stripped profile should be approximately equal to that of the unstripped profile within the tidal radius.
@@
The $J$-factor and structural properties of subhaloes are also known to be related. For instance, [70] examined the relationship between $J$ and $G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max}$, where $v_{\rm max}$ and $r_{\rm max}$ are the maximum circular velocity ($v_{c}=\sqrt{GM(<r)/r}$), and corresponding radius. They show that while for an NFW profile $\log_{10}J=1.23\log_{10}(G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max})$, their tidally-stripped simulations of NFW profiles have a slightly shallower slope of 0.86. Our NFWT model agrees remarkably well with their findings, but differences in the profile model can cause the $J$-factor to vary by a factor of $\sim 2$.
@@
Figure 10: Relationship between the $J$-factor, and the combination of structural parameters $G^{-1}v^{4}_{\rm max}r^{-1}_{\rm max}$, where $v_{\rm max}$ and $r_{\rm max}$ are the maximum circular velocity. The slope of 1.23 is what is expected for an NFW profile, while 0.86 is what has been found for tidally stripped NFW profiles [70]. We find that our NFWT model agrees with this scaling relation, but differences in initial profiles can cause a deviation in the $J$-factor by factors of $\approx 2$.
@@ 4.4 Model sensitivity in annihilation: Summary
In this section, we considered the effects of tidal mass loss on the inner structure of subhaloes as they relate to calculations of the annihilation rate by calculating quantities related to the dark matter annihilation rate for NFW, EinCore, and EinCusp profiles. By using a model assuming a truncation of the distribution function in energy space, we are able to examine a scenario where the central density is conserved during tidal stripping across different initial profile models. We compared our prediction for NFW profiles to two widely used approximations; first, a model in which the original NFW profile is sharply truncated such that it contains the same bound mass (ST model), and secondly, the parameterization from [16] that was tuned to simulation results (H03). The ST model will overestimate the density within the tidal radius by construction. By contrast, the Hayashi profile is known to underestimate the central density.
@@
Comparing first the energy-truncated NFW and Einasto profiles, we found that for fixed mass, the concentration, flux multiplier (measured within the tidal radius), and $J$-factor of the EinCuspT profile is typically 10-50 per cent higher than the NFWT profile, while the EinCoreT profile is typically lower. These differences reflect the different central densities of the initial models, as the annihilation signal is dominated by these high-density regions.
@@
Then, comparing tidal stripping models (the NFWT model versus the commonly used ST and H03 approximations), we found that neither of the common approximations agrees well with the Energy Truncation model. The effective concentration at fixed mass is lower by a factor of 10 for both approximations. Both models under-predict the flux multiplier within the tidal radius by a factor of $\sim 100$. However, this result is mainly due to the differences in the tidal radii, and if we consider the evolution of $f_{M}/f_{M,{\rm init}}$ the H03 and ST model underestimate and overestimate the flux multiplier by $\sim 20$ per cent, respectively. If we consider the $f_{M}/f_{M,{\rm un}}$; i.e., the flux multiplier from our model compared to a truncated profile within the same radius, $r_{\rm t}$, we can find that the two approximations underestimate the flux multiplier by an order of magnitude. Overall, the $J$-factor is expected to be overestimated in the ST model and underestimated in the H03 model; the differences between these models are a factor of 5 once the halo has been stripped to $\sim 5\%$ of its initial mass.
@@
In summary, subhalo concentrations, flux multipliers , and $J$-factors of tidally stripped haloes are sensitive to the innermost part of the density profile, and predicted fluxes can vary by a factor of 5 or more depending on profile assumptions. Thus, it is important to have physically motivated models that allow us to predict the annihilation of subhaloes down to arbitrarily small bound mass fractions that are not resolved in simulations. The Energy Truncation Model allows us to explore the tidal evolution of the central density of subhaloes under the assumption that cuspy centres are preserved in tidal stripping. Our results suggest that current constraints on the CDM annihilation cross-section may be inaccurate due to their underlying assumptions about subhalo density profiles; addressing these inaccuracies will be the focus of future work.
@@ 4.5 Implications for the Milky Way
We provide a brief exploration of what our results mean in the context of the Milky Way galaxy—more robust calculations will be the focus of future work. Throughout this section, we assume that the Milky Way has a virial mass of $M_{\rm MW}=10^{12}M_{\odot}$, a virial radius of 250 kpc, and a concentration of 10. We use a Planck 2018 cosmology [71].
We assume that the Milky Way is created from the mergers of haloes that follow the distribution ${\rm d}^{2}N/{\rm d}M_{i}{\rm d}z$, where $N$ is the number of haloes, $M_{i}$ is the mass of a subhalo at infall redshift, $z$. We calculate this distribution by assuming the infall population is the same as the field population:
$\dfrac{{\rm d}^{2}N}{{\rm d}M_{i}{\rm d}z}=\dfrac{1}{\rho_{c}(z)}\dfrac{{\rm d}M_{\rm MW}(z)}{{\rm d}z}\dfrac{{\rm d}n}{{\rm d}M}\,\,\,,$ (4.5)
@@
where ${\rm d}n(M,z)/{\rm d}M$ is the halo mass function from [72], $M_{\rm MW}(z)$ is the mass of the Milky Way at redshift $z$, and $\rho_{c}$ is the critical density. Effectively, this specifies the volume of the Universe that must collapse to assemble a halo of mass $M_{\rm MW}$. Finally, we assume the Milky Way mass evolves according to [52]:
$M_{\rm MW}(z)=M_{\rm MW}\exp(-8.2z/c)\,\,\,.$ (4.6)
As the Milky Way is formed, some of the substructure will persist as self-bound haloes, with a present day mass of $M_{0}$. The rest of the dark matter will contribute to the smooth component of the halo. This smooth component can be calculated from:
$M_{\rm smooth}=\int_{\infty}^{0}\int_{M_{\rm min}}^{M_{\rm MW}}\dfrac{{\rm d}N}{{\rm d}M_{i}{\rm d}z}(M_{i}-M_{0}){\rm d}M_{i}{\rm d}z\,\,\,.$ (4.7)
We assume a minimum halo mass of $10^{-6}\,M_{\odot}$, as in [2].
We model the relationship between the infall mass and the current mass of the subhaloes using [73]:
$M_{0}=M_{i}\exp\left(\dfrac{-0.81\sqrt{200}z_{i}}{\pi}\right)\,\,\,.$ (4.8)
We note that this parameterization does not take into account the effect of subhalo mass profile on mass-loss rates, and is likely an overestimate of the tidal mass loss.
Then, the total Milky Way annihilation signal can be expressed as
$J_{\rm MW}=J_{\rm sub}+J_{\rm smooth}\,\,\,.$ (4.9)
where $J_{\rm sub}$ is the annihilation signal due to Milky Way subhaloes at redshift $z=0$ and $J_{\rm smooth}$ is the annihilation signal from the smooth halo component.
The signal from the smooth component, $J_{\rm smooth}$, is calculated using the mass from Equation (4.7), and assuming an NFW profile. The signal from the substructure can be found by adding the contribution from each subhalo,
$J_{\rm sub}=\int_{\infty}^{0}\int_{M_{\rm min}}^{M_{\rm MW}}\dfrac{{\rm d}N}{{\rm d}M_{i}{\rm d}z}J(M_{0}){\rm d}M_{i}{\rm d}z\,\,\,,$ (4.10)
The annihilation signal of each subhalo, $J=J(M_{0})$, depends on the assumed density profile of the subhalo, the subhalo mass, and concentration. We use different models for the subhalo density profiles and assign concentrations to subhaloes, assuming they lie on the median concentration–mass relation from [74].
@@
The top panel of Figure 11 shows the relative contribution of the substructure to the total annihilation signal as a function of subhalo mass. The signal is dominated by the smallest subhaloes, with approximately 70 per cent of the signal coming from haloes less than $10^{6}\,M_{\odot}$. The contribution of subhaloes to the total annihilation signal differs slightly between models, with small subhaloes in more centrally concentrated models contributing slightly more to the annihilation signal.
@@
Then, we compared the ‘substructure boost factor’ [2]—defined as the annihilation flux from substructure, $J_{\rm sub}$ relative to the flux from the smooth component of the halo—in the bottom panel Fig 11. We calculated the smooth component by assuming all of the Milky Way mass was smoothly distributed in an NFW profile with mass [^6] We find $\log({M_{\rm tot}/M_{\odot}})$ = 11.3 which is slightly lower than our specified $\log({M_{\rm MW}/M_{\odot}})$, but agrees to within an order of magnitude.
@@
We find that including substructure increases the signal by a factor of $\sim 50$ for the more centrally concentrated profiles (ST and EinCuspT), and $\sim 10$ times for the less centrally concentrated profiles (H03). The lower values are comparable to the predictions of [2], who explored different concentration–mass relations, but found a substructure boost factor of $\sim$16 when assuming a H03 model and concentration relations similar to ours [^7] We note that our quantity is slightly different—we are showing the contribution from substructure relative to a smooth signal calculated using the total Milky Way Mass, while [2] uses the smooth signal from the mass not contained in subhaloes.. This demonstrates the overall importance of substructure in calculating the annihilation signal in the Milky Way, but also the large range of values predicted depending on the assumed subhalo density profile. Assuming low-mass subhaloes have profiles closest to the EinCuspT or NFWT models, we predict a ‘substructure boost factor’ of as much as 20–50, relative to the contribution from the smooth halo.
We note, however, that there are several simplifications in this analysis. First, our parameterization for subhalo stripping (Equation 4.8) overestimates how quickly profiles are disrupted, and does not take into account major mergers. Second, we assumed that all haloes lie on the median concentration–mass relation and that the Milky Way has built up its mass at an average rate. Finally, the rate at which haloes are stripped is dependent on the subhalo model, with cuspier profiles experiencing less stripping—this, however, would increase the difference between the models. We leave a full investigation of how the assumption that cuspy centres are preserved in tidal stripping affects the annihilation signal for future work.